OpenStudy (anonymous):

For f(x)= 1/x-5 and g(x)= x^2 + 2 , find (f o g)(x) and find (g o f) (6)? I really need help getting these answers.

OpenStudy (ranga):

(f o g)(x) is same as f(g(x)). You need to find f(g(x)) for x = 6. So what is f(g(6))? g(x) = x^2 + 2. Find g(6) ----- Step 1 f(x) = 1/x - 5 f(g(6)) = Substitute g(6) found in Step 1 for x in the equation in the previous line.

OpenStudy (anonymous):

So the answer would be 1/33?

OpenStudy (ranga):

Your answer is correct IF \[f(x) = \frac{ 1 }{ (x-5) }\] But not correct IF \[f(x) = \frac{ 1 }{ x } - 5\]

OpenStudy (anonymous):

It was the first one, so I guess the answer is correct. Thanks! I'm still not sure on the answer to the (f o g)(x) though?

OpenStudy (ranga):

Another way to do it: \[f(x) = \frac{ 1 }{ (x - 5) }\] \[g(x) = x ^{2} + 2\] \[(f o g)(x) = \frac{ 1 }{ ((x ^{2} + 2) - 5) } = \frac{ 1 }{ (x ^{2} - 3) }\] \[(fog)(6) = \frac{ 1 }{ (6^{2} - 3) } = \frac{ 1 }{ 33 }\]

OpenStudy (ranga):

Do the same for (g o f)(x) for x = 6

OpenStudy (anonymous):

Sorry I'm confused right now, was the 1/33 answer for (f o g)(x) or for (g o f)(6)

OpenStudy (ranga):

(f o g)(x) = 1/33 for x = 6

OpenStudy (anonymous):

Oh, so for (f o g)(x) they just want me to pick a random number for x and solve and we picked 6?

OpenStudy (anonymous):

So if 1/33 is the answer to the first question what is (g o f)(6)

OpenStudy (ranga):

I assumed they meant x = 6 for both parts. But if not you can leave it as (f o g)(x) = 1/(x^2 - 3) But you can evaluate (g o f)(x) for x = 6. But you will have to find (g o f)(x) first. Use the same technique we used earlier.

OpenStudy (anonymous):

So the answer would be 3? because 1/6-5 = 1, then 1^2 + 2= 3

OpenStudy (ranga):

Yes it is 3. But don't forget to use parenthesis 1/6 - 5 above should be 1/(6-5) = 1.

OpenStudy (anonymous):

Ok, thank you! This is a study guide to a big test this week, so I have a few more questions haha. But thanks for the help on this one!

OpenStudy (ranga):

So I would leave the answers as: (f o g)(x) = 1/(x^2 - 3) (g o f)(6) = 3

OpenStudy (ranga):

Sure no problem. Good luck on the tset.

OpenStudy (anonymous):

That's what I did, thanks!